OpenStudy (aaaa):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution. One point for each.

OpenStudy (anonymous):

There are two equations to this. The first equation is 2x + 8 = -6 Do you know the second equation?

OpenStudy (aaaa):

its only one

OpenStudy (anonymous):

Not quite, compound means two, the second equation is 2x + 8 = 6 Please notice that one has a negative sign and the other does not

OpenStudy (anonymous):

I will solve the first equation 2x + 8 = -6 by subtracting 8 from both sides 2x = -14 and dividing both sides by 2 x = -7 Can you solve the second equation?